OpenStudy (anonymous):

Ratios The small rectangle had length of x and width of x + 3 The large rectangle had a length of 3x and a width of x^2 + 7x + 12

OpenStudy (anonymous):

I know I'd end up having: 3(x^2 + 7x + 12) = 3x(x + 3)

OpenStudy (anonymous):

which would be: 3x^2 + 21x + 24 = 3x^2 + 9x

OpenStudy (anonymous):

is x^3 + 4x^2 + 3x but what is the simplified form?

OpenStudy (anonymous):

let me know if i did anything wrong in the steps

OpenStudy (anonymous):

HOW can i put it in simplified form?

OpenStudy (mertsj):

\frac{\[\frac{3}{3x}=\frac{x+3}{(x+4)(x+3)}\]

OpenStudy (mertsj):

\[\frac{1}{x}=\frac{1}{x+4}\]

OpenStudy (mertsj):

\[x+4=x\]

OpenStudy (mertsj):

\[4=0\]

OpenStudy (mertsj):

No solution

OpenStudy (anonymous):

But these are my options: 3(x + 4) x + 3 /x^2 + 7x + 12 3x /x + 4 1 /3x + 4

OpenStudy (mertsj):

I don't think you posted the problem correctly because you ask what is the value of x in the proportion and then you list a bunch of answers that contain x.

OpenStudy (mertsj):

So either you have posted the wrong question or the wrong proportion.

OpenStudy (anonymous):

Oh Im so sorry!!! I posted the wrong question! Its the simplified proportion I think